Shift Summary for owl shift, 12/16/99

John Shields (jshields@ksera)
Thu, 16 Dec 1999 07:10:25 -0600

Shift Summary for owl shift, 12/16/99

On Shift: J. Shields, S. TaegAr

Hours of beam available at all intensities: 0.0
Hours of beam available at nominal intensity: 0.0
Hours of E799 data (special runs not included): 0.0
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time:

Run II beam studies
========================================================================
Runs ended this shift:

15230 -- COSMICS

Tapes brought to FCC this shift:
(including tapes for online split)

========================================================================
Accesses during shift: 1

AG fiddled with KAMI equipment

========================================================================
Overview of shift:

ran Cosmics, labeled tapes. Numerous ktevana errors
and pipeline problems.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

15230 01:30 after multiple failed attempts to fix laser
illumination missing quadrant by resetting
pipeline, end run and do drt_fresh_start

15231 01:40 beginning new cosmics run. Laser illumination
is fixed.

04:05 JS, AG go on access so that AG can fiddle with
the KAMI apparatus.

04:20 return from access.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

do a drt_fresh_start before beginning the next run

Ar/Eth West will need to be changed soon. Also, all ktev flashlights/
flashlight batteries are dead.